Write an equation of a line that passes through (-12, -14) with slope 6.

Answer by mananth(16946)

m= 6

Plug value of the slope and point(-12,-14) in
y=mx+b
-14.00 = 6 * -12 +b
b= -14 + 72

b= 58
So the equation will be
Y = 6 x + 58
m.ananth@hotmail.ca